# CADSpotting: Robust Panoptic Symbol Spotting on Large-Scale CAD Drawings

### v3

@@ Abstract
We introduce CADSpotting, an effective method for panoptic symbol spotting in large-scale architectural CAD drawings. Existing approaches struggle with symbol diversity, scale variations, and overlapping elements in CAD designs. CADSpotting overcomes these challenges by representing primitives through densely sampled points with attributes like coordinates and colors, using a unified 3D point cloud model for robust feature learning. To enable accurate segmentation in large, complex drawings, we further propose a novel Sliding Window Aggregation (SWA) technique, combining weighted voting and Non-Maximum Suppression (NMS). Moreover, we introduce LS-CAD, a new large-scale CAD dataset to support our experiments, with each floorplan covering around 1,000 m2, significantly larger than previous benchmarks. Experiments on FloorPlanCAD and LS-CAD datasets show that CADSpotting significantly outperforms existing methods. We also demonstrate its practical value through automating parametric 3D reconstruction, enabling interior modeling directly from raw CAD inputs.
1 Introduction
Architectural Computer-Aided Design (CAD) drawings, especially those used for interior design, serve as detailed digital representations that convey essential information about a building’s structure, layout, and intricate details. These drawings include critical elements such as furniture placement, electrical layouts, and spatial organization, ensuring consistency and precision throughout the construction process. For instance, the Shanghai Mercedes-Benz Arena utilize over 3,000 CAD drawings to guide its structural design, optimize spatial arrangements, and streamline the construction workflow. As CAD designs continue to grow in complexity, automating the detection and interpolation of symbols within these floorplans becomes increasingly important in various downstream tasks, such as code compliance checking and 3D interior modeling [20].
@@ 1 Introduction
Traditional methods approach symbol recognition in a query-by-example manner [29], but these techniques struggle to handle the vast diversity of graphical symbols in real-world datasets. More recent learning-based approaches [25, 7] have shown significant progress in addressing the symbol spotting task. Early solutions [5, 27, 40] typically leverage convolutional neural networks (CNNs) or graph-based models to spot symbols. More advanced methods leverage Transformer architectures and attention mechanisms [6, 40], improving global relationship reasoning among symbols. These methods use rasterized pixel images derived from vector CAD drawings as input. Such conversion, however, introduces discrepancies between the original vector graphics and the rasterized images, resulting in errors due to the loss of precise geometric information.
The representative work SymPoint [16] treats CAD drawings as sets of primary points, leveraging techniques from point cloud analysis for effective CAD symbol spotting. Such a representation simplifies model complexity while achieving strong performance on panoptic symbol spotting tasks in CAD drawings. Building on this foundation, SymPoint-V2 [17] introduces Layer feature-enhanced encoding to incorporate graphical layer information and a position-guided Training method to enhance model learning and accelerate convergence. Yet both approaches rely on manually defined primitive feature representations, i.e., four fixed primitive types. As a result, they struggle to capture the full diversity of CAD graphical primitives. In addition, these methods focus on processing small to medium-scale datasets due to the absence of effective strategies for preserving connectivity during symbol spotting across expansive drawings. By far, there is still a lack of feasible methods to address the panoptic symbol spotting task in large-scale CAD drawings, as dense symbol clutter and significant variations in scale present substantial challenges for accurate segmentation and symbol recognition.
@@
In this paper, we present CADSpotting, a simple yet highly effective panoptic symbol spotting method tailored for handling CAD drawings at a much larger scale. Specifically, CADSpotting does not rely on fixed graphic primitive types. Instead, it densely samples points along CAD graphic primitives to construct comprehensive point cloud representations. Each point is represented by basic attributes, such as coordinates and color, forming a lightweight but expressive feature set. We utilize a unified 3D point cloud processing model to learn robust features from the sampled points. CADSpotting then employs a streamlined Transformer decoder for panoptic symbol spotting, effectively leveraging the learned features to achieve precise segmentation. To handle large-scale, real-world CAD drawings, we further propose a novel Sliding Window Aggregation (SWA) technique, combining a weighted voting strategy with Non-Maximum Suppression (NMS) to enable accurate and effective panoptic segmentation.
@@
On the data front, we further introduce LS-CAD, a new large-scale CAD dataset containing 50 annotated floorplans from a variety of building types, including campuses and office complexes. Each drawing follows the same fine-grained annotation standards as the FloorPlanCAD dataset, with an average coverage of over 1,000 square meters per floorplan. LS-CAD is the first dataset of its scale in this domain, part of which will be released to the research community under a license waiver, to stimulate further developments. Experimental results on both FloorPlanCAD and LS-CAD demonstrate that CADSpotting outperforms the state-of-the-art methods by a margin, demonstrating its robustness and effectiveness in handling the complexities of large-scale, real-world CAD drawings.
We also demonstrate the effectiveness of CADSpotting by automating the generation of 3D interior models through parametric modeling. Using precise instance-level data obtained from CADSpotting, we compute spatial parameters for architectural elements, including wall contours, door and window positions, door orientations, and pivot points. With these spatial parameters clearly defined, we leverage parametric reconstruction techniques within Blender to efficiently build 3D interior models directly from raw CAD data. This approach significantly streamlines the modeling process, enhancing productivity and accuracy. To the best of our knowledge, our proposed pipeline is the first comprehensive solution that seamlessly integrates CAD segmentation with automated 3D interior model reconstruction.
2 Related Work
@@
Image segmentation is a fundamental task in computer vision, traditionally divided into two main categories: semantic segmentation [18, 2, 34, 1, 36] and instance segmentation [28, 8, 9, 13]. Semantic segmentation assigns a class label to every pixel in an image, while instance segmentation distinguishes individual objects within the same class. Recently, SAM [13] introduces a large model trained on over one billion masks, leveraging prompt-based interactions for enhanced interactive segmentation. SAM2 [24] extends this approach to video segmentation. However, semantic and instance segmentation methods struggle to accurately represent uncountable background elements, such as the sky or terrain. To address this challenge, Kirillov et al. [12] propose Panoptic Segmentation, which unifies semantic and instance segmentation, effectively handling both countable and uncountable background components. Early approaches [11, 31, 15, 3] primarily employ CNN-bases architectures, while recent advancements, such as Mask2Former [4], SegFormer [36], and OneFormer [9], leverage Transformer models, incorporating a mask attention mechanism to boost segmentation performance. Following the large-model paradigm of SAM, SEEM [41] further refines segmentation using a prompt-driven design. Despite these advances, most existing segmentation models remain pixel-centric, limiting their effectiveness in processing vector graphics like CAD drawings. Consequently, they often fail to capture critical vector-based information, such as overlapping relationships and precise geometric attributes.
@@
Early efforts [22, 23, 26, 10, 37] in symbol spotting on CAD drawings rely on handcrafted feature descriptors, such as shape and structure, coupled with techniques like sliding window searches or graph matching for symbol retrieval. The introduction of deep learning techniques marks a significant shift, with models based on Faster R-CNN [25] and YOLO v3 [7] greatly enhancing the accuracy of symbol recognition tasks [27]. However, traditional symbol spotting methods are primarily designed for countable objects and do not effectively handle uncountable symbols, which are common in CAD environments. To address this limitation, Fan et al. [5] introduce the FloorPlanCAD dataset and CNN-GCN method, enabling the recognition of both countable and uncountable symbols for more comprehensive CAD parsing. Countable object instances (e.g., windows, doors) can be clearly identified and individually counted whereas uncountable stuff (e.g., wall, railing) refers to elements that cannot be individually distinguished. Building on this, Zheng et al. [40] propose GAT-CADNet, a model that represents CAD drawings as graph structures and performs segmentation by predicting adjacency matrices. Further advancements include CADTransformer [6], which utilizes the Vision Transformer to extract features from scalar maps, leading to improved predictions. Meanwhile, significant progress has been made in the field of 3D point cloud segmentation [39, 35, 14, 30]. The seminal work SymPoint [16] proposes a novel method that converts symbols into primary point representations, thereby enhancing feature extraction through point cloud analysis and upsampling techniques. Concurrently, Liu et al. [17] extend SymPoint by incorporating LFE and PGT modules, further improving feature representation and model performance. Despite these advances, the manually defined feature representation is often constrained by a limited set of graphical primitive types, which hinders their ability to capture the full diversity of CAD symbols.
3 Method
@@
3.1 Primitive Feature Learning
@@
SymPoint [16] introduces a new concept that represents CAD primitives as a set of 2D points. It constructs a handcrafted feature for each graphical primitive, capturing details such as type (e.g., line, arc, circle, ellipse), length, etc. SymPoint then applies point cloud analysis to learn primitive representations, further enhancing connectivity through an Attention with Connection Module. A masked-attention transformer decoder is used to complete the panoptic symbol spotting task. Although SymPoint shows strong performance in CAD symbol spotting, its feature representation is manually defined and limited to four fixed primitive types, which may constrain its capacity to capture the diverse range of CAD symbols. Real-world CAD drawings often contain various complex elements, such as Bézier curves, rather than simple predefined line types. Therefore, methods that rely on handcrafted features lack sufficient robustness. To address these constraints, we propose a novel graphical primitive feature representation for CAD drawings that incorporates dense point sampling within a unified point cloud analysis model to improve generalizability of primitive-wise representations.
Dense Point Sampling.
In CAD drawings, graphical primitives are typically categorized into line types and shape types. Line primitives encompass straight segments, curves, and complex contours such as Bézier curves, while shape primitives include basic geometric forms like rectangles, circles, and ellipses. Inspired by SymPoint [16], we propose a dense point sampling strategy that performs equidistant sampling on each primitive to generate dense point data for feature extraction. Rather than assigning fixed types to CAD graphical primitives and using handcrafted features to represent them, our dense point sampling method offers greater feasibility and adaptability, making it suitable for a wide range of CAD primitives. Fig.3 illustrates the process of our dense point sampling method.
Given a CAD drawing containing primitives, we perform dense sampling to generate a 2D point cloud from these primitives at a fixed distance . Each 2D point is represented by a six-dimensional vector: ,where represent the 2D coordinates of the point, is set to 0, and represent the color information (red, green, blue channels) of the point. We then employ Point Transformer V3 [35] (PTv3) as the backbone for feature extraction, leveraging its design optimized for the unordered nature of point clouds, which provides robust performance in feature extraction.
Feature Pooling.
@@
To obtain the final feature representation for CAD primitives, we apply Primitive Mixed Pooling (PMP) to transform point-wise features into primitive-wise features. Let denote our pooling operator, which aggregates the point features into primitive-wise feature vector :
@@
where denotes the total number of sampled points, and represents the number of primitives. Let represent the single primitive with its corresponding dense point cloud . The -th dimension feature of is determined by the sum of the maximum and average values of the corresponding -th dimension features across all points:
@@
This pooling method condenses a large set of dense point features into a more compact set of primitive-level features, effectively reducing computational complexity and storage requirements. At the same time, primitive-level pooling retains the essential information of the contained points, preserving global consistency in the primitive representation and minimizing information loss. Loss calculations are also conducted at the primitive-level, aligning with our final goal of performing CAD spotting tasks.
3.2 Panoptic Symbol Spotting
@@
We utilize a simple transformer decoder to extract semantic and instance symbol spotting information from the learned primitive features. In our model, the pooled primitive features serve as key and value inputs to three consecutive layers of Transformer decoders, where the self-attention mechanism effectively captures dependencies among the input features. The queries are initialized randomly and optimized during training to compute accurate self-attention scores with other features.
@@
Following OneFormer3D [14], we partition the decoder output into two parts: the first represent instance proposals, and the subsequent outputs correspond to semantic symbol spotting. The output dimension of the + vector is , integrating both instance and semantic information. The first dimensions are used to generate instance masks, while the remaining dimensions determine the semantic category of each primitive. This configuration enables panoptic symbol spotting by integrating instance and semantic information for comprehensive scene understanding.
@@
We translate decoder outputs to semantic proposals and instance proposals. A semantic proposal includes the log-likelihood values for each primitive being classified into different semantic labels and can be further processed through thresholding to generate a binary mask. Similarly, an instance proposal also contains log-likelihood values, but these indicate the likelihood of each primitive being classified as a specific instance. Additionally, we employ a bipartite matching strategy based on the Hungarian algorithm to find the optimal correspondence between predicted results and ground truth labels while training. During inference, we select the top- instance proposals of the highest confidence scores and apply matrix-NMS [33] to suppress non-maximal instance predictions.
@@
Loss Function. The loss function is composed of classification loss and primitive mask loss. The classification loss is defined as a multi-class cross-entropy loss. The primitive mask loss combines binary cross-entropy loss, , with the Dice loss [21], , to evaluate the correspondence between predicted instances and ground truth.
3.3 Sliding Window Aggregation
@@
To address the challenge of panoptic symbol segmentation at large scales, a key issue is ensuring the generalization of instance symbol segmentation across varying scales. During inference on larger drawings than those encountered during training, the network must produce significantly more instance proposals. However, this process is limited by the fixed number of instance queries inherent in Mask Transformer-based methods, a constraint that also affects our approach. Nonetheless, CAD instance segmentation offers certain advantages over general image instance segmentation. Specifically, in CAD floorplans, instances of the same category maintain relatively consistent sizes across different scenes when a uniform scale is applied. Consequently, we can reasonably assume that the size of instances in inference will not exceed that of the same category in the training images.
Therefore, we propose using a fixed-size window based on the training image dimensions and apply 2D sliding window inference on the large input drawings. For each sliding window, once the primitives intersects with or are encompassed by the window, we utilize the features of these primitives along with the window’s semantic and instance proposals, thereby increasing the likelihood of fully capturing potential instances. After gathering proposals across all windows, we apply distinct aggregation strategies for semantic and instance proposals.
For semantic proposals, we use a voting scheme to predict each primitive’s semantic label, treating each primitive as an independent unit. Primitives that extend across larger spatial areas may be captured by multiple windows. In such cases, we prioritize proposals that capture the primitive as completely as possible by using weighted voting, where each window’s voting weight is determined by the proportion of the primitive’s dense points observed relative to the total dense points of that primitive.
@@
For instance proposals, we utilize Sparse Non-Maximum Suppression(Sparse-NMS) on the proposals generated from all windows, which increases processing efficiency (detailed analysis is provided in the supplementary material). A potential instance may be partially or fully observed across multiple windows. By ensuring a sufficiently small sliding step size, each instance is fully captured within at least one window. Incomplete proposals from other windows are subsequently filtered out during the NMS process.
3.4 Automated 3D Interior Reconstruction
@@
CADSpotting delivers precise panoptic segmentation results, serving as the basis for our parametric 3D interior modeling algorithm. Utilizing detailed instance information and primitive positions from segmentation, we apply instance-level processing techniques to compute spatial parameters for key architectural elements, specifically walls, doors, and windows. For doors, we determine their positions, orientations, and pivot points by analyzing the relationships between arcs and lines within each segmented instance. Window positions, on the other hand, are directly extracted from their corresponding instances. Complex wall structures require a specialized approach: we first traverse and merge endpoints of adjacent lines within wall instances, forming closed polygons. These polygons are then rasterized into binary images, from which final wall contours are identified using the method from [32]. With these spatial parameters, we reconstruct a complete 3D interior model using a parametric reconstruction pipeline in Blender. The necessary 3D primitives for walls, doors, and windows are manually prepared in Blender, providing reusable templates for streamlined parametric interior modeling. This approach significantly enhances efficiency, automating the conversion of raw CAD data into structurally accurate 3D interior models within minutes through procedural generation. An example result from our automated 3D interior reconstruction method is illustrated in Fig. 4, with additional details provided in the supplementary material.
@@
4 Experiments
4.1 The Proposed LS-CAD Dataset
We introduce LS-CAD, a new large-scale CAD dataset consisting of 50 floorplans from extensive buildings, such as campuses and office buildings. Each CAD drawing in LS-CAD covers an area of at least 1,000 square meters, with the number of primitives ranging from approximately 2,900 to over ten thousand. Table.1 showcases 3 large-scale CAD drawings from our dataset, providing detailed information such as area, number of primitives and instance count, which underscore the large-scale nature of LS-CAD.
Additionally, the floorplans in LS-CAD include a wide variety of complex symbols, such as doors, windows, walls, elevators and parking spaces. We provide fine-grained annotations consistent with the standards of the FloorPlanCAD [5] dataset. A portion of LS-CAD will be publicly available under a license waiver to support further research in panoptic symbol spotting for large-scale CAD drawings.
4.2 Experimental Settings
Implementation Details. The FloorPlanCAD dataset comprises approximately 15,000 drawings, annotated across 35 symbol categories, including 30 things and 5 stuff classes. To ensure fair comparison, we utilize the same training subset from the FloorPlanCAD dataset as used by SymPoint. We train our model on a machine with 8 NVIDIA A100 GPUs, using a batch size of 2 per GPU for 512 epochs. For dense point sampling, we set a fixed distance . We use the AdamW [19] optimizer with a learning rate of and a weight decay of 0.05. Data augmentation techniques include random horizontal flipping with a probability of , global rotation and scaling transformations, translation along the x and y axes and color normalization to . We set loss weight as . Additionally, we set as the number of semantic labels, and both the top- value and are set to 220. We use as the step size of SWA on LS-CAD dataset.
Metrics. Following the approach of [6], we assess the performance of our model using multiple metrics. For semantic symbol spotting, we use F1 score, weighted F1 score (wF1), and mean Intersection over Union (mIoU). For instance symbol spotting, we employ AP50, AP75, and mean Average Precision (mAP). For panoptic symbol spotting, we utilize Panoptic Quality (PQ), Segmentation Quality (SQ), and Recognition Quality (RQ). PQ is a comprehensive metric combining SQ and RQ to evaluate model performance in semantic and instance segmentation tasks. Further details on metric formulation are available in [6].
4.3 Quantitative Evaluation
| Methods | AP50 | AP75 | mAP |
| DINO [38] | 64.0 | 54.9 | 47.5 |
| SymPoint [16] | 66.3 | 55.7 | 52.8 |
| CADSpotting (ours) | 72.2 | 69.1 | 69.0 |
@@ 4.3 Quantitative Evaluation
| PanCADNet [5] | 59.5 | 82.6 | 66.9 | 65.6 | 86.1 | 76.1 | 58.7 | 81.3 | 72.2 |
| CADTransformer [6] | 68.9 | 88.3 | 73.3 | 78.5 | 94.0 | 83.5 | 58.6 | 81.9 | 71.5 |
| GAT-CADNet [40] | 73.7 | 91.4 | 80.7 | - | - | - | - | - | - |
| SymPoint [16] | 83.3 | 91.4 | 91.1 | 84.1 | 94.7 | 88.8 | 48.2 | 69.5 | 69.4 |
| CADSpotting (Ours) | 88.9 | 95.6 | 93.0 | 89.7 | 96.2 | 93.2 | 80.6 | 89.7 | 89.8 |

### v4

@@ Abstract
We introduce CADSpotting, an effective method for panoptic symbol spotting in large-scale architectural CAD drawings. Existing approaches often struggle with symbol diversity, scale variations, and overlapping elements in CAD designs, and typically rely on additional features (e.g., primitive types or graphical layers) to improve performance. CADSpotting overcomes these challenges by representing primitives through densely sampled points with only coordinate attributes, using a unified 3D point cloud model for robust feature learning. To enable accurate segmentation in large drawings, we further propose a novel Sliding Window Aggregation (SWA) technique that combines weighted voting and Non-Maximum Suppression (NMS). Moreover, we introduce LS-CAD, a new large-scale dataset comprising 45 finely annotated floorplans, each covering approximately 1,000 , significantly larger than prior benchmarks. LS-CAD will be publicly released to support future research. Experiments on FloorPlanCAD and LS-CAD demonstrate that CADSpotting significantly outperforms existing methods. We also showcase its practical value by enabling automated parametric 3D interior reconstruction directly from raw CAD inputs.
Introduction
Architectural Computer-Aided Design (CAD) drawings, especially those used for interior design, serve as detailed digital representations that convey essential information about a building’s structure, layout, and intricate details. These drawings include critical elements such as furniture placement, electrical layouts, and spatial organization, ensuring consistency and precision throughout the construction process. For instance, the Shanghai Mercedes-Benz Arena utilizes over 3,000 CAD drawings to guide its structural design, optimize spatial arrangements, and streamline the construction workflow. As CAD designs continue to grow in complexity, automating the detection and interpolation of symbols within these floorplans becomes increasingly important in various downstream tasks, such as code compliance checking and 3D interior modeling (MarsBIM 2024).
@@ Introduction
Traditional query-by-example based methods (Rusiñol, Lladós, and Sánchez 2010) struggle with the diverse and complex graphical symbols in real-world CAD drawings. Learning-based approaches (Ren et al. 2016; Farhadi and Redmon 2018) have advanced the field, with early methods using convolutional neural networks (CNNs) or graph-based models (Fan et al. 2021; Rezvanifar, Cote, and Albu 2020; Zheng et al. 2022), and later works incorporating Transformers and attention mechanisms (Fan et al. 2022; Zheng et al. 2022) for improved global reasoning. However, these approaches typically operate on rasterized images derived from vector CAD data, leading to loss of geometric fidelity and recognition errors. SymPoint (LIU et al. 2024) avoids rasterization by treating CAD drawings as point sets and applying point cloud techniques for effective symbol spotting. SymPoint-V2 (Liu et al. 2024) builds on this by introducing layer-aware encoding and position-guided training. While effective, both methods rely on predefined primitive types, and SymPoint-V2 additionally depends on graphical layer information, limiting their ability to generalize across diverse real CAD data. Moreover, they focus on small to medium-scale drawings and lack mechanisms for preserving spatial consistency in large-scale drawings. As a result, panoptic symbol spotting in expansive CAD drawings remains challenging due to dense clutter, scale variation, and segmentation complexity.
@@
In this paper, we present CADSpotting, a simple yet effective panoptic symbol spotting method tailored for handling CAD drawings at a much larger scale. Unlike prior approaches, CADSpotting does not rely on fixed graphic primitive types or additional layer information. Instead, it densely samples points along CAD graphic primitives to construct a compact yet expressive point cloud representation, where each point is represented solely by its coordinate attributes. A unified 3D point cloud model is employed to learn robust features from these sampled points, followed by a streamlined Transformer decoder that performs panoptic symbol spotting. To handle large-scale, real-world CAD drawings, we introduce a novel Sliding Window Aggregation (SWA) technique that combines weighted voting and Non-Maximum Suppression (NMS) to enable accurate and scalable panoptic segmentation across windowed regions.
@@
On the data front, we further introduce LS-CAD, a new large-scale CAD dataset containing 45 finely annotated floorplans from diverse building types, including campuses and office complexes. Each floorplan spans over 1,000 on average and follows the same fine-grained annotation standards as FloorPlanCAD. LS-CAD is the first dataset of its scale in this domain, and it will be released under a license waiver to support future research. Experimental results on both FloorPlanCAD and LS-CAD demonstrate that CADSpotting outperforms the state-of-the-art methods, demonstrating its robustness and effectiveness in real-world scenarios. Finally, we demonstrate CADSpotting’s practical value by automating 3D interior reconstruction. Leveraging instance-level outputs, we extract spatial parameters such as wall contours, door/window positions, orientations, and pivot points. These parameters are then used to drive parametric modeling in Blender, enabling efficient and accurate 3D interior generation directly from raw CAD inputs. To our knowledge, this constitutes the first end-to-end pipeline that tightly integrates CAD symbol spotting with parametric 3D reconstruction (Fig.1).
Related Work
@@
Image segmentation is a fundamental task in computer vision, traditionally divided into two main categories: semantic segmentation (Long, Shelhamer, and Darrell 2015; Chen et al. 2017; Wu et al. 2019; Chen 2017; Xie et al. 2021) and instance segmentation (Ronneberger, Fischer, and Brox 2015; He et al. 2017; Jain et al. 2023; Kirillov et al. 2023). While semantic segmentation labels each pixel by class, instance segmentation further distinguishes between object instances of the same class. Recently, SAM (Kirillov et al. 2023) introduces a large model trained on over one billion masks, leveraging prompt-based interactions for enhanced interactive segmentation. SAM2 (Ravi et al. 2024) extends this approach to video segmentation. However, these methods often struggle to represent uncountable background elements (e.g., sky, terrain). To unify countable and uncountable regions, Kirillov et al. (Kirillov et al. 2019) introduce panoptic segmentation, which unifies semantic and instance segmentation, effectively handling both countable and uncountable background components. Early approaches(Sun et al. 2019; Li et al. 2019; Chen, Wang, and Qiao 2020) use CNN backbones, while recent advances like Mask2Former (Cheng et al. 2022) and OneFormer (Jain et al. 2023) leverage Transformer architectures with mask attention mechanisms for improved accuracy. Large models such as SEEM (Zou et al. 2024) further refines segmentation using a prompt-driven design. Despite these advances, most methods remain pixel-centric, limiting their effectiveness in processing vector graphics like CAD drawings. As a result, they struggle to capture critical vector-based information, including precise geometry and overlapping relationships.
@@
Early CAD symbol spotting methods (Nguyen, Tabbone, and Terrades 2008; Nguyen, Tabbone, and Boucher 2009; Rezvanifar, Cote, and Branzan Albu 2019; Jiang et al. 2021; Yang et al. 2023) rely on handcrafted feature descriptors, such as shape and structure, coupled with techniques like sliding window searches or graph matching for symbol retrieval. The introduction of deep learning significantly improved performance, with models based on Faster R-CNN (Ren et al. 2016) and YOLO (Farhadi and Redmon 2018) achieving higher accuracy (Rezvanifar, Cote, and Albu 2020), though they primarily focus on countable object instances. To address uncountable elements, Fan et al.(Fan et al. 2021) introduce the FloorPlanCAD dataset and a CNN-GCN framework for holistic parsing. Zheng et al.(Zheng et al. 2022) further model CAD drawings as graphs, using adjacency prediction for segmentation. CADTransformer (Fan et al. 2022) adopt Vision Transformers to extract features from scalar maps. Inspired by advancements in 3D point cloud segmentation (Zhao et al. 2021; Wu et al. 2024; Kolodiazhnyi et al. 2024; Schult et al. 2023), SymPoint (LIU et al. 2024) reformulates CAD drawings as primary point sets and applies point-based learning for effective symbol spotting. SymPoint-V2 (Liu et al. 2024) extends this with layer-aware encoding and position-guided training. Despite their effectiveness, both methods rely on a fixed set of primitive types, and SymPoint-V2 assumes the availability of reliable layer information. These assumptions often do not hold in real-world CAD data, limiting their ability to capture the full diversity and complexity of CAD symbols.
Method
@@
Primitive Feature Learning
@@
SymPoint (LIU et al. 2024) represents CAD primitives as sets of 2D points and constructs handcrafted features for each, encoding attributes such as type (line, arc, circle, ellipse) and length. It then applies point cloud analysis with an Attention with Connection Module to learn primitive representations, followed by a masked-attention Transformer decoder for panoptic symbol spotting. While effective, SymPoint is limited by its reliance on manually defined features and a fixed set of four primitive types, which restricts its ability to handle the full diversity of CAD symbols. Real-world drawings often include complex elements like Bézier curves, making handcrafted features insufficiently robust. To overcome these limitations, we propose a novel primitive representation based on dense point sampling and unified point cloud learning, enabling more generalizable and expressive primitive-level features.
Dense Point Sampling
In CAD drawings, graphical primitives are generally classified into line types (e.g., straight segments, curves, Bézier contours) and shape types (e.g., rectangles, circles, ellipses). Inspired by SymPoint (LIU et al. 2024), we adopt a dense point sampling strategy that performs equidistant sampling along each primitive to generate dense point-level data for feature learning. Unlike methods that rely on fixed primitive types and handcrafted features, our sampling approach is more flexible and generalizable, making it suitable for a wide variety of CAD elements. Fig. 3 illustrates the sampling process. Given a CAD drawing with primitives, we sample points at a fixed interval to construct a 2D point cloud , where each point is represented as a 3D vector with . We then use Point Transformer V3 (Wu et al. 2024) (PTv3) as the backbone to extract robust features, leveraging its design optimized for unordered point cloud data.
Feature Pooling
@@
To obtain the final feature representation for CAD primitives, we use Primitive Mixed Pooling (PMP) to convert point-wise features into primitive-wise features .
@@
where is the total number of sampled points, and is the number of primitives. Let denote the dense point set corresponding to the -th primitive. The -th dimension of the primitive-level feature is computed as the sum of the maximum and average values of the -th dimension across all points in :
@@
This pooling strategy compresses dense point features into a compact set of primitive-level representations, reducing computational and storage overhead. Meanwhile, it preserves essential geometric information, maintaining global consistency and minimizing information loss. Loss computation is also performed at the primitive level, aligning with the objective of CAD symbol spotting.
Panoptic Symbol Spotting
@@
We utilize a lightweight Transformer decoder to extract semantic and instance-level information from the pooled primitive features. These features serve as keys and values for three consecutive decoder layers, where self-attention captures dependencies among primitives. Query embeddings are randomly initialized and optimized during training to compute accurate self-attention scores with other features.
@@
Following OneFormer3D (Kolodiazhnyi et al. 2024), the decoder output is divided into two parts: the first vectors correspond to instance proposals, and the remaining vectors represent semantic symbol spotting. The output has shape , where the first dimensions are used to predict instance masks, and the rest indicate semantic class scores. This unified representation enables panoptic symbol spotting by combining both semantic and instance-level predictions.
@@
Each semantic proposal includes log-likelihoods for each primitive being classified into different semantic labels and can be further processed through thresholding to generate a binary mask. Similarly, an instance proposal also contains log-likelihood values, and these indicate the likelihood of each primitive being classified as a specific instance. During training, we apply bipartite matching using the Hungarian algorithm to align predictions with ground truth. At inference, we select the top- instance proposals with the highest confidence and apply matrix-NMS (Wang et al. 2020) to eliminate redundant predictions.
@@
Loss Function The overall loss function consists of a classification loss and a primitive mask loss. The classification loss is defined as a multi-class cross-entropy loss. The mask loss combines binary cross-entropy and Dice loss (Milletari, Navab, and Ahmadi 2016) , jointly measuring the alignment between predicted instance masks and ground truth.
Sliding Window Aggregation
@@
To address the challenge of panoptic symbol segmentation at large scales, a key issue is generalizing instance segmentation across drawings of varying sizes. During inference, larger inputs require significantly more instance proposals than those seen during training. However, Mask Transformer-based methods, including ours, are constrained by a fixed number of instance queries. Fortunately, CAD floorplans offer an advantage: instances of the same category typically maintain consistent sizes across different scenes under uniform scaling. This allows us to assume that instance sizes at inference do not exceed those in training data.
Based on this assumption, we adopt a fixed-size sliding window strategy, using windows aligned with training dimensions to process large input drawings. For each window, we collect all primitives it covers or intersects and apply the decoder to generate local semantic and instance proposals. Aggregated results across all windows are then merged using distinct strategies for semantic and instance outputs.
For semantic proposals, we apply a weighted voting scheme to assign each primitive a class label. When a primitive appears in multiple windows, we prioritize proposals from windows that observe a larger portion of it. The voting weight is proportional to the number of observed dense points relative to the total points of the primitive.
@@
For instance proposals, we use Sparse Non-Maximum Suppression (Sparse-NMS) to merge results from overlapping windows efficiently (see supplementary material). By using a sufficiently small step size, each instance is fully captured in at least one window. Incomplete detections from other windows are eliminated during the NMS process.
Automated 3D Interior Reconstruction
@@
CADSpotting produces accurate panoptic segmentation results that serve as the foundation for our parametric 3D interior modeling pipeline. Using instance-level segmentation and primitive positions, we extract spatial parameters for key architectural elements, including walls, doors, and windows. For doors, we compute positions, orientations, and pivot points by analyzing geometric relationships between arcs and lines within each segmented instance. Window positions are directly derived from their corresponding instances. Wall modeling involves a more specialized process: we merge endpoints of adjacent lines within wall instances to form closed polygons, which are rasterized into binary masks. Final wall contours are then extracted using the method from (Suzuki and be 1985). These spatial parameters are fed into a parametric reconstruction pipeline implemented in Blender. Predefined 3D assets for walls, doors, and windows are used as reusable templates, enabling efficient and automated generation of structurally accurate 3D interiors. This approach converts raw CAD inputs into detailed 3D models within minutes. An example result is shown in Fig. 4, with further details provided in the supplementary material.
@@
Experiments
The Proposed LS-CAD Dataset
We introduce LS-CAD, a new large-scale CAD dataset comprising 45 floorplans from expansive buildings, such as campuses and office complexes. Each drawing covers at least 1,000 square meters, with the number of primitives ranging from approximately 2,900 to over 10,000. Table. 1 presents three representative examples, detailing area, primitive count, and instance count to highlight the large-scale nature of LS-CAD. The dataset features a rich variety of complex symbols, including doors, windows, walls, elevators, and parking spaces. All floorplans are annotated with fine-grained labels consistent with the FloorPlanCAD (Fan et al. 2021) standard. The LS-CAD dataset will be publicly released to support future research in panoptic symbol spotting for large-scale CAD drawings.
Experimental Settings
Implementation Details
The FloorPlanCAD dataset comprises approximately 15,000 drawings, annotated across 35 symbol categories, including 30 things and 5 stuff classes. To ensure fair comparison, we utilize the same training subset from the FloorPlanCAD dataset as used by SymPoint. We train our model on a machine with 8 NVIDIA A100 GPUs, using a batch size of 2 per GPU for 512 epochs. For dense point sampling, we set a fixed distance . We use the AdamW (Loshchilov 2017) optimizer with a learning rate of and a weight decay of 0.05. Data augmentation techniques include random horizontal flipping with a probability of , global rotation and scaling transformations, translation along the x and y axes. We set loss weight as . Additionally, we set as the number of semantic labels, and both the top- value and are set to 220. We use as the step size of SWA on LS-CAD dataset.
Metrics
Following the approach of (Fan et al. 2022), we assess the performance of our model using multiple metrics. For semantic symbol spotting, we use F1 score and weighted F1 score (wF1). For instance symbol spotting, we employ AP50, AP75, and mean Average Precision (mAP). For panoptic symbol spotting, we utilize Panoptic Quality (PQ), Segmentation Quality (SQ), and Recognition Quality (RQ). PQ is a metric combining SQ and RQ to evaluate model performance in semantic and instance segmentation tasks. Further details on metric formulation are available in (Fan et al. 2022).
Quantitative Evaluation
| Method | AP50 | AP75 | mAP |
| DINO (Zhang et al. 2022) | 64.0 | 54.9 | 47.5 |
| SymPoint (LIU et al. 2024) | 66.3 | 55.7 | 52.8 |
| CADSpotting (ours) | 70.6 | 66.6 | 66.3 |
@@
| PanCADNet (Fan et al. 2021) | 59.5 | 82.6 | 66.9 | 65.6 | 86.1 | 76.1 | 58.7 | 81.3 | 72.2 |
| CADTransformer (Fan et al. 2022) | 68.9 | 88.3 | 73.3 | 78.5 | 94.0 | 83.5 | 58.6 | 81.9 | 71.5 |
| GAT-CADNet (Zheng et al. 2022) | 73.7 | 91.4 | 80.7 | - | - | - | - | - | - |
| SymPoint (LIU et al. 2024) | 83.3 | 91.4 | 91.1 | 84.1 | 94.7 | 88.8 | 48.2 | 69.5 | 69.4 |
| CADSpotting (Ours) | 87.4 | 93.5 | 93.4 | 88.3 | 94.2 | 93.7 | 71.5 | 82.9 | 86.3 |